OpenStudy (andu1854):

Find the derivative of: tan(x^2+x)^5

OpenStudy (thadyoung):

You need to apply the chain rule for this. First take the derivative of the outside function multiplied by the derivative of the inner function. So, here you have: \(\sf \color{blue}{[tan(x^2+x)]^5} \) which is almost the same thing as thinking: \(\sf \color{red}{a^5}\), where a = tan(x\(^2\)+x) So you have the first derivative of the outer function is: \(\sf \color{green}{5[tan(x^2+x)]^4} \times\frac{d}{dx}tan(x^2+x)\) where you can see it as: \(\sf \frac{d}{dx}\color{orange}{tan(b)}\) where b = x\(^2\)+x then do the same thing with the LAST inner function, and multiply by that again.

OpenStudy (andu1854):

thank you very much... I see where I made my error...